OpenStudy (anonymous):

plzz find the area under the curve of function 3^(2x+1)^0.5 from x=0 to x=4

OpenStudy (anonymous):

integral

OpenStudy (anonymous):

I understand it's \[3^{\sqrt{2x+1}}\] Substitute \[2x+1 =y^2\] Differentiate both sides and you will get a equation in terms of y. Adjust the limits accordingly and integrate. I believe you will get \[2\int\limits_{?}^{?}3^{y}ydy\] Resubstitute on finding the integral.

OpenStudy (anonymous):

\[\int\limits_0^4 3^{\sqrt{1+2 x}} \, dx=\frac{-24+78 \text{Log}[3]}{\text{Log}[3]^2}=51.1138 \]Plot attached.